Re: Shift report 20030114 00:00-08:00

From: Peter H. L. Christiansen (pchristi@nbi.dk)
Date: Fri Jan 17 2003 - 11:13:09 EST

  • Next message: Apache: "Shift report 20030117 08:00-16:00"
    Hi
    
    Looking a bit more careful at run 7374 I found that the fibres were all on 
    and seems to work fine (at least for T1). I will finish the first 
    calibrations early next week of T1 and T2. I am still interested in the 
    z-positions for TPM2 and TPM1 used last year.
    
    Sorry for the inconvenience
        Peter
    
    On Wed, 15 Jan 2003, Peter H. L. Christiansen wrote:
    
    > Hi
    > 
    > > Horizontal fibers in FS were turned on this morning at 08:15. They have
    > > been off until now, but were sadly missed in Copenhagen.
    > > 
    > 
    > Looking at run 7374 I could not see that anything changed i.e., all the 
    > adc values are 0. 
    > 
    > Do we need the fibres ? I think it is a very fast way to estimate the
    > drift velocity in all the TPCs, so it would be good if someone could look
    > into this (or decide that we don't do it!). If we decide to do it I could
    > include the calibrations for TPM1 and TPM2 in the script if the person
    > that did this last year using BrTpcDriftFibCalModule could send the lines
    > for TPM1 and TPM2. BTW I have not checked if the TPM2 fibre data is there.
    > 
    > Cheers
    >    Peter
    > 
    > -- :-) --------------------------- )-:
    >  Peter H L Christiansen @ NBI
    >  EMAIL  : pchristi@nbi.dk
    >  OFFICE : Tb1@NBI  (353 25269)
    >  HOME   : Vesterbrogade 204, 2. tv 
    >  PHONE  : 40840492(mob.) 
    > :-D --------------------------- \-:
    > 
    > 
    
    -- 
    :-) --------------------------- )-:
     Peter H L Christiansen @ NBI
     EMAIL  : pchristi@nbi.dk
     OFFICE : Tb1@NBI  (353 25269)
     HOME   : Vesterbrogade 204, 2. tv 
     PHONE  : 40840492(mob.) 
    :-D --------------------------- \-:
    


    This archive was generated by hypermail 2.1.5 : Fri Jan 17 2003 - 11:13:37 EST